# Empowering Targeted Neighborhood Search via Hyper Tour for Large-Scale TSP

### v2

@@ 1 Introduction
The traveling salesman problem (TSP) is a well-known NP-hard combinatorial optimization problem. Given a list of cities (or vertices) and the distances between each pair of cities, TSP aims to find the shortest possible route that visits each city (or vertex) exactly once and returns to the origin city. Due to its complexity and broad application value in areas such as content delivery (lee2022economics), robot routing (macharet2018survey), biology (lenstra1975some), circuit design (ducomman2016alternative), Web crawling (tripathi2021age) and service scheduling (kiraly2015redesign), the TSP has consistently drawn significant attention from both industry and research communities (li2024fast; grinsztajn2023winner; chalumeau2024combinatorial; luo2024neural; kwon2020pomo; jin2023pointerformer; drakulic2023bq; lyu2024scalable; zhou2023towards; son2023meta; li2025unify; wang2024asp; xiao2024distilling; li2023distribution; min2024unsupervised; ma2023learning; ye2023deepaco; wang2025efficient; li2025unify).
Recent advances in TSP research have increasingly focused on addressing the significant challenges of large-scale instances, where the primary difficulties lie in high memory consumption and the exponentially growing search space as the problem size increases. Traditional methods for large-scale TSPs, such as LKH3 (helsgaun2017extension), typically employ heuristic guidance to identify promising candidate vertices during the search, thereby reducing the search space and memory consumption. However, they rely heavily on numerous expert-crafted rules and iterative searches, making them inflexible and time-consuming, especially for large-scale TSP instances. More recently, with the introduction of neural networks into TSP solving (vinyals2015pointer), researchers have begun exploring or combining various neural models with traditional methods to enhance TSP performance on large-scale TSPs, which can be roughly divided into heatmap-guided [^1] Heatmaps are typically generated by GCNs, where each element represents the probability of an edge being part of the optimal TSP solution. methods (fu2021generalize; qiu2022dimes; sun2023difusco; xia2024position; luoboosting) and decomposition-based methods (fang2024invit; ye2024glop; cheng2023select; zhou2025dualopt; zhengudc).
@@
To address these challenges, we propose a novel framework that integrates heatmap-guided and decomposition-based approaches while mitigating their respective limitations for efficiently solving large-scale TSPs. To this end, we introduce the Hyper Tour Guided Neighborhood Search framework (HyperNS), which systematically addresses memory bottlenecks, lack of global guidance, and poor initialization, enabling efficient routing in large-scale instances. Motivated by the idea of “clustering first, route second(gillett1974heuristic), our method first constructs a sparse heatmap graph using a sub-graph sampling-then-aggregate paradigm, significantly reducing the memory footprint compared with full $\mathbb{R}^{n\times n}$ representations while retaining crucial edge information. Large-scale instances are then partitioned and abstracted into supernodes by iteratively deleting bridge edges, and a hyper tour is generated over these supernodes by solving a reduced problem. This hyper tour serves as a high-level abstraction of the global path, providing consistent guidance for both initialization and subsequent optimization. Guided by the hyper tour, we perform tour initialization to generate high-quality initial solutions, followed by a targeted neighborhood search that iteratively refines inter-supernode connections and reorganizes intra-supernode structures, focusing exclusively on relevant edges and their neighbors. By strategically combining sparse representation, hyper tour guidance, and targeted neighborhood search, HyperNS narrows the search space, preserves global coherence, mitigates memory bottlenecks, and guides the optimization process toward faster convergence and higher-quality solutions. Experimental results on both generated and real-world datasets, spanning various scales of TSP instances, demonstrate the effectiveness of our method compared with existing neural methods. The main contributions of this work are as follows:
@@ 2 Related Work
Early neural methods, typically limited to instances with up to 1,000 vertices, can be divided into constructive-based methods (vinyals2015pointer; bello2016neural; deudon2018learning; kool2018attention; khalil2017learning; li2024fast; grinsztajn2023winner; chalumeau2024combinatorial; luo2024neural; kwon2020pomo; jin2023pointerformer; drakulic2023bq; lyu2024scalable; jacobs21reinforcement; zhang2022learning; jiang2022learning; bi2022learning; zhou2023towards; son2023meta; li2025unify; wang2024asp), which generate tours from scratch, and improvement-based methods (joshi2019efficient; xiao2024distilling; li2023distribution; min2024unsupervised; chen2019learning; wu2021learning; lu2019learning; hudson2022graph; yao22reversible; ma2023learning; ye2023deepaco; wang2025efficient; li2025unify), which first generate a solution and then iteratively refine it under neural guidance toward better solutions. Most neural methods for large-scale TSP instances follow the improvement-based paradigm, with one exception INViT (fang2024invit), which suffers from much lower performance. In this section, we focus on improvement-based methods for large-scale TSPs, which can be broadly divided into heatmap-guided methods and decomposition-based methods depending on how neural guidance is applied, while other methods are reviewed in Appendix A.1.
@@
Heatmap-guided methods provide global guidance by predicting a probability distribution over all edges. AttGCRN+MCTS (fu2021generalize) extends the Att-GCRN (graph convolutional residual network with attention mechanism) model to generate heatmaps and uses MCTS for optimization, handling up to 10,000 vertices. DIMES (qiu2022dimes), DIFUSCO (sun2023difusco) and SoftDist (xia2024position) improve upon this with meta-learning, diffusion models and distance based heatmap generation, respectively, but struggle with generalization to larger instances (¿10,000) due to MCTS’s memory constraints.
@@
Decomposition-based methods reduce the problem scale by dividing a large instance into one or multiple subproblems, which are solved independently, thereby alleviating memory issues. Representative examples include GLOP (ye2024glop), Select&Optimize (cheng2023select), DualOpt (zhou2025dualopt), UDC (zhengudc), and PRC (luoboosting). While these approaches are more scalable, their lack of global guidance makes them prone to inefficient search and local optima. Different from these approaches, DRHG (li2025destroy) adopts a destroy-and-repair framework, which iteratively removes certain path nodes and compresses the remaining contiguous edge segments into hyper-edges, thereby significantly reducing the problem-solving scale. However, its destroy phase relies on predefined rules without a well-justified rationale and lacks explicit global guidance, leading to excessive iterations that slow down problem solving. As a result, it is prone to getting trapped in local optima and performs even worse than earlier methods on large-scale instances. In contrast, our method introduces a higher-level hyper tour abstraction that serves as a proxy for the global path, providing consistent global guidance for both initialization and iterative refinement, thereby overcoming DRHG’s local bias.
@@
Different from most existing neural methods, we design an efficient routing method for large-scale TSP instances, capable of handling up to 70K vertices, by leveraging a cluster-based hyper tour as global guidance. Unlike heatmap-based approaches that rely on storing multiple large matrices, our method uses sparse heatmap graph generation combined with hyper tour-guided targeted neighborhood search, which significantly reduces memory usage while maintaining strong performance on large-scale TSPs. Compared with divide-and-conquer based methods (e.g., GLOP (ye2024glop), Select&Optimize (cheng2023select), DualOpt (zhou2025dualopt), UDC (zhengudc), and PRC (luoboosting)), which refine subtours independently without global coordination, HyperNS constructs a global hyper tour that guides both initialization and optimization. In particular, our initialization could speed up convergence and improve overall results, while targeted neighborhood search enforces global consistency by reorganizing connections between clusters and refining local structures. This synergy among improved initialization, global abstraction, and local refinement clearly distinguishes HyperNS from existing methods and underpins its superior performance.
@@ 3.2 Model Overview
In this section, we briefly introduce our HyperNS for efficient routing using hyper tours for larger-scale TSPs. Fig. 1 provides an overview of our HyperNS. First, considering that heatmap can help identify local structures by highlighting high-probability edges, we enhance Att-GCRN by incorporating geometric information, generate sub-heatmaps via edge-covering-based subgraph sampling, and merge them into a sparse heatmap graph by selecting top-k elements. We then partition this graph into clusters, each abstracted as a supernode, and generate a hyper tour by constructing a reduced TSP over these supernodes and solving it using Att-GCRN-GI and LK search (DBLP:journals/ior/LinK73)  [^2] The Lin–Kernighan (LK) search is currently one of the most efficient and effective algorithms for small-scale TSP instances. It is widely used as the basic search unit in methods such as LKH and MCTS due to its strong performance. For more details, please refer to Appendix A.5.. The hyper tour serves as an abstraction of the global TSP solution at a higher level, providing consistent guidance for both initialization and subsequent optimization. Using the hyper tour as a guide, we propose a tour initialization procedure, followed by a targeted neighborhood search process. By restricting the optimization to hyper-tour-related edges, our method reduces the search space and progressively refines both supernode connections and intra-supernode links, ultimately yielding better solutions.
@@ 3.3 Sparse Heatmap Graph Construction
Considering that the heatmap generated by neural networks like GCN (khalil2017learning) and Att-GCRN (joshi2019efficient) is adopted by recent methods for large-scale TSPs, we aim to use the heatmaps as guiding signals to create a hyper tour at higher level for subsequent optimization. However, the current sub-graph sampling-then-aggregation paradigm (fu2021generalize; qiu2022dimes; sun2023difusco) has three issues, including missing geometry information like distances and triangle areas among vertices, random sampling without considering the coverage of uncovered vertices and the evaluation of all edges, and high memory usage for using full heatmaps. These issues not only hinder the model’s ability to generalize to larger-scale problems but also reduce its solving efficiency and effectiveness.
To address these challenges, we first propose a self-supervised pretraining task inspired by (zhou2023improving) to capture geometric information like edge lengths and triangle areas and enhance the vertex representations of Att-GCRN, which is referred to as Att-GCRN-GI. Second, we introduce an edge-covering-oriented graph sampling method that models graph sampling as a set-covering problem (young2008greedy), aiming to minimize the number of subgraphs and sampling time while ensuring comprehensive edge coverage and then get a sparse heatmap graph directly from the sub-heatmaps. We next detail the construction of the sparse heatmap graph with Att-GCRN-GI in the following, while leaving the architectural details of Att-GCRN-GI to Appendix A.2, as it is not the primary focus of this work.
Specifically, for each vertex $v_{i}$ of $\mathbb{G}$, we first get its corresponding subgraph $\mathbb{G}_{i}=(\mathcal{N}(v_{i},p),\mathcal{E}(\mathcal{N}(v_{i},p)))$, where $\mathcal{N}(v_{i},p)$ is to get $v_{i}$’s $p$ nearest neighbors and $\mathcal{E}(S)=\{(u,v)\in\mathbb{E}|u,v\in S\}$. We regard $C_{i}=\mathcal{E}(\mathcal{N}(v_{i},p))$ as a candidate set and can obtain subgraphs by identifying candidate sets that cover all edges for $n$ sets, thereby addressing a set-covering problem. Next, we apply a greedy algorithm (young2008greedy) to solve it, which iteratively selects the set containing the largest number of uncovered edges at each stage.
@@ 3.4 Hyper Tour Generation and Initialization
Motivated by the idea of “clustering first, route second” (gillett1974heuristic), we propose addressing large-scale TSP instances by dividing them into clusters based on sparse heatmap graphs, and then pursuing a hyper tour among these clusters. Complex optimal solutions for large-scale TSP instances often exhibit smaller, well-organized local structures, such as subloops or clusters. These local structures are crucial to the overall solution, as they capture the optimal arrangement and connections of adjacent or related elements. Heatmaps can help identify these local structures by highlighting edges with high local probabilities, facilitating the grouping of vertices within the same sub-tour into a cluster. Each cluster is then abstracted into a supernode, a single vertex representing all vertices within the cluster—allowing us to form a smaller TSP instance while preserving the essential structure of the original graph. We solve this reduced problem and refer to its solution as the “hyper tour”, which determines how the supernodes are connected and serves as a guide for generating and optimizing the initial solution.
@@ 3.5 Targeted Neighborhood Search
Table 1: Performance on generated datasets (luoboosting). “N/A” indicates failure to solve due to memory constraints.
@@
Concorde (cook2011traveling) 23.12 (0%) 5.78h N/A N/A N/A N/A N/A N/A N/A N/A
LKH3 (helsgaun2017extension) 23.12 (0%) 1.69h 50.90 (0%) 2.66h 71.77 (0%) 2.52h 101.31 (0%) 7.33h 159.93 (0%) 50.3h
AM-greedy (kool2018attention) 33.55 (45.1%) 3.45m 96.74 (90.1%) 4.03m 153.42 (113.8%) 2.13m 244.20 (141.0%) 5.13m N/A N/A
GCN-BS (joshi2019efficient) 51.1 (121.0%) 3.23h N/A N/A N/A N/A N/A N/A N/A N/A
AttGCRN+MCTS (fu2021generalize)
@@
DIMES (qiu2022dimes) 23.69 (2.5%) 4.62h 52.37 (2.9%) 3.47h 74.06 (3.2%) 3.57h N/A N/A N/A N/A
DIFUSCO (sun2023difusco) 23.39 (1.2%) 0.41h 52.04 (2.2%) 0.71h 73.62 (2.6%) 0.79h N/A N/A N/A N/A
SoftDist (xia2024position) 23.63 (2.2%) 9.88m 52.25 (2.7%) 1.03h 74.03 (3.1%) 1.05h N/A N/A N/A N/A
SO (cheng2023select) 23.77 (2.8%) 0.93h 52.60 (3.3%) 1.92h 74.30 (3.5%) 2.03h 105.03 (3.7%) 8.92h N/A N/A
UDC (zhengudc) 23.53(1.8%) 0.57h 33.26(2.5%) 0.21h N/A N/A N/A N/A N/A N/A
DRHG (li2025destroy) 23.19 (0.3%) 2.31h 51.39 (0.9%) 3.13h 72.85 (1.3%) 4.26h 103.75 (2.4%) 6.14h 167.23 (4.6%) 9.35h
INViT (fang2024invit) 24.50 (6.0%) 0.26h 54.19 (6.5%) 0.76h 76.12 (6.1%) 0.73h 108.49 (7.1%) 5.68h 171.38 (7.2%) 16.76h
PRC (luoboosting) 23.37 (1.1%) 3.47h 51.67 (1.3%) 6.23h 73.08 (1.8%) 3.71h 103.66 (2.3%) 5.93h 163.95 (2.5%) 16.92h
DualOpt (zhou2025dualopt) 23.31 (0.8%) 0.31h 51.56 (1.3%) 0.64h 72.62 (1.2%) 0.87h 102.90 (1.6%) 2.28h 162.81 (1.8%) 7.83h

### v3

@@ 1 Introduction
The traveling salesman problem (TSP) is a well-known NP-hard combinatorial optimization problem. Given a list of cities (or vertices) and the distances between each pair of cities, TSP aims to find the shortest possible route that visits each city (or vertex) exactly once and returns to the origin city. Due to its complexity and broad application value in areas such as content delivery (Lee et al., 2022), robot routing (Macharet and Campos, 2018), biology (Lenstra and Kan, 1975), circuit design (Ducomman et al., 2016), Web crawling (Tripathi et al., 2021) and service scheduling (Király and Abonyi, 2015), the TSP has consistently drawn significant attention from both industry and research communities (Li et al., 2024; Grinsztajn et al., 2023; Chalumeau et al., 2023; Luo et al., 2024; Kwon et al., 2020; Jin et al., 2023; Drakulic et al., 2023; Lyu et al., 2024; Zhou et al., 2023a; Son et al., 2023; Li et al., 2025b; Wang et al., 2024; Xiao et al., 2024; Li et al., 2023; Min et al., 2024; Ma et al., 2023; Ye et al., 2023; Wang et al., 2025; Li et al., 2025b).
Recent advances in TSP research have increasingly focused on addressing the significant challenges of large-scale instances, where the primary difficulties lie in high memory consumption and the exponentially growing search space as the problem size increases. Traditional methods for large-scale TSPs, such as LKH3 (Helsgaun, 2017), typically employ heuristic guidance to identify promising candidate vertices during the search, thereby reducing the search space and memory consumption. However, they rely heavily on numerous expert-crafted rules and iterative searches, making them inflexible and time-consuming, especially for large-scale TSP instances. More recently, with the introduction of neural networks into TSP solving (Vinyals et al., 2015), researchers have begun exploring or combining various neural models with traditional methods to enhance TSP performance on large-scale TSPs, which can be roughly divided into heatmap-guided [^1] Heatmaps are typically generated by GCNs, where each element represents the probability of an edge being part of the optimal TSP solution. methods (Fu et al., 2021; Qiu et al., 2022; Sun and Yang, 2023; Xia et al., 2024; Luo et al., 2025) and decomposition-based methods (Fang et al., 2024; Ye et al., 2024; Cheng et al., 2023; Zhou et al., 2025; Zheng et al., 2024).
@@
To address these challenges, we propose a novel framework that integrates heatmap-guided and decomposition-based approaches while mitigating their respective limitations for efficiently solving large-scale TSPs. To this end, we introduce the Hyper Tour Guided Neighborhood Search framework (HyperNS), which systematically addresses memory bottlenecks, lack of global guidance, and poor initialization, enabling efficient routing in large-scale instances. Motivated by the idea of “clustering first, route second(Gillett and Miller, 1974), our method first constructs a sparse heatmap graph using a sub-graph sampling-then-aggregate paradigm, significantly reducing the memory footprint compared with full $\mathbb{R}^{n\times n}$ representations while retaining crucial edge information. Large-scale instances are then partitioned and abstracted into supernodes by iteratively deleting bridge edges, and a hyper tour is generated over these supernodes by solving a reduced problem. This hyper tour serves as a high-level abstraction of the global path, providing consistent guidance for both initialization and subsequent optimization. Guided by the hyper tour, we perform tour initialization to generate high-quality initial solutions, followed by a targeted neighborhood search that iteratively refines inter-supernode connections and reorganizes intra-supernode structures, focusing exclusively on relevant edges and their neighbors. By strategically combining sparse representation, hyper tour guidance, and targeted neighborhood search, HyperNS narrows the search space, preserves global coherence, mitigates memory bottlenecks, and guides the optimization process toward faster convergence and higher-quality solutions. Experimental results on both generated and real-world datasets, spanning various scales of TSP instances, demonstrate the effectiveness of our method compared with existing neural methods. The main contributions of this work are as follows:
@@ 2 Related Work
Early neural methods, typically limited to instances with up to 1,000 vertices, can be divided into constructive-based methods (Vinyals et al., 2015; Bello et al., 2017; Deudon et al., 2018; Kool et al., 2019; Khalil et al., 2017; Li et al., 2024; Grinsztajn et al., 2023; Chalumeau et al., 2023; Luo et al., 2024; Kwon et al., 2020; Jin et al., 2023; Drakulic et al., 2023; Lyu et al., 2024; Jacobs et al., 2021; Zhang et al., 2022; Jiang et al., 2022; Bi et al., 2022; Zhou et al., 2023a; Son et al., 2023; Li et al., 2025b; Wang et al., 2024), which generate tours from scratch, and improvement-based methods (Joshi et al., 2019; Xiao et al., 2024; Li et al., 2023; Min et al., 2024; Chen and Tian, 2019; Wu et al., 2021a; Lu et al., 2019; Hudson et al., 2022; Yao et al., 2022; Ma et al., 2023; Ye et al., 2023; Wang et al., 2025; Li et al., 2025b), which first generate a solution and then iteratively refine it under neural guidance toward better solutions. Most neural methods for large-scale TSP instances follow the improvement-based paradigm, with one exception INViT (Fang et al., 2024), which suffers from much lower performance. In this section, we focus on improvement-based methods for large-scale TSPs, which can be broadly divided into heatmap-guided methods and decomposition-based methods depending on how neural guidance is applied, while other methods are reviewed in Appendix A.1.
@@
Heatmap-guided methods provide global guidance by predicting a probability distribution over all edges. AttGCRN+MCTS (Fu et al., 2021) extends the Att-GCRN (graph convolutional residual network with attention mechanism) model to generate heatmaps and uses MCTS for optimization, handling up to 10,000 vertices. DIMES (Qiu et al., 2022), DIFUSCO (Sun and Yang, 2023) and SoftDist (Xia et al., 2024) improve upon this with meta-learning, diffusion models and distance based heatmap generation, respectively, but struggle with generalization to larger instances (¿10,000) due to MCTS’s memory constraints.
@@
Decomposition-based methods reduce the problem scale by dividing a large instance into one or multiple subproblems, which are solved independently, thereby alleviating memory issues. Representative examples include GLOP (Ye et al., 2024), Select&Optimize (Cheng et al., 2023), DualOpt (Zhou et al., 2025), UDC (Zheng et al., 2024), and PRC (Luo et al., 2025). While these approaches are more scalable, their lack of global guidance makes them prone to inefficient search and local optima. Different from these approaches, DRHG (Li et al., 2025a) adopts a destroy-and-repair framework, which iteratively removes certain path nodes and compresses the remaining contiguous edge segments into hyper-edges, thereby significantly reducing the problem-solving scale. However, its destroy phase relies on predefined rules without a well-justified rationale and lacks explicit global guidance, leading to excessive iterations that slow down problem solving. As a result, it is prone to getting trapped in local optima and performs even worse than earlier methods on large-scale instances. In contrast, our method introduces a higher-level hyper tour abstraction that serves as a proxy for the global path, providing consistent global guidance for both initialization and iterative refinement, thereby overcoming DRHG’s local bias.
@@
Different from most existing neural methods, we design an efficient routing method for large-scale TSP instances, capable of handling up to 70K vertices, by leveraging a cluster-based hyper tour as global guidance. Unlike heatmap-based approaches that rely on storing multiple large matrices, our method uses sparse heatmap graph generation combined with hyper tour-guided targeted neighborhood search, which significantly reduces memory usage while maintaining strong performance on large-scale TSPs. Compared with divide-and-conquer based methods (e.g., GLOP (Ye et al., 2024), Select&Optimize (Cheng et al., 2023), DualOpt (Zhou et al., 2025), UDC (Zheng et al., 2024), and PRC (Luo et al., 2025)), which refine subtours independently without global coordination, HyperNS constructs a global hyper tour that guides both initialization and optimization. In particular, our initialization could speed up convergence and improve overall results, while targeted neighborhood search enforces global consistency by reorganizing connections between clusters and refining local structures. This synergy among improved initialization, global abstraction, and local refinement clearly distinguishes HyperNS from existing methods and underpins its superior performance.
@@ 3.2 Model Overview
In this section, we briefly introduce our HyperNS for efficient routing using hyper tours for larger-scale TSPs. Fig. 1 provides an overview of our HyperNS. First, considering that heatmap can help identify local structures by highlighting high-probability edges, we enhance Att-GCRN by incorporating geometric information, generate sub-heatmaps via edge-covering-based subgraph sampling, and merge them into a sparse heatmap graph by selecting top-k elements. We then partition this graph into clusters, each abstracted as a supernode, and generate a hyper tour by constructing a reduced TSP over these supernodes and solving it using Att-GCRN-GI and LK search (Lin and Kernighan, 1973)  [^2] The Lin–Kernighan (LK) search is currently one of the most efficient and effective algorithms for small-scale TSP instances. It is widely used as the basic search unit in methods such as LKH and MCTS due to its strong performance. For more details, please refer to Appendix A.5.. The hyper tour serves as an abstraction of the global TSP solution at a higher level, providing consistent guidance for both initialization and subsequent optimization. Using the hyper tour as a guide, we propose a tour initialization procedure, followed by a targeted neighborhood search process. By restricting the optimization to hyper-tour-related edges, our method reduces the search space and progressively refines both supernode connections and intra-supernode links, ultimately yielding better solutions.
@@ 3.3 Sparse Heatmap Graph Construction
Considering that the heatmap generated by neural networks like GCN (Khalil et al., 2017) and Att-GCRN (Joshi et al., 2019) is adopted by recent methods for large-scale TSPs, we aim to use the heatmaps as guiding signals to create a hyper tour at higher level for subsequent optimization. However, the current sub-graph sampling-then-aggregation paradigm (Fu et al., 2021; Qiu et al., 2022; Sun and Yang, 2023) has three issues, including missing geometry information like distances and triangle areas among vertices, random sampling without considering the coverage of uncovered vertices and the evaluation of all edges, and high memory usage for using full heatmaps. These issues not only hinder the model’s ability to generalize to larger-scale problems but also reduce its solving efficiency and effectiveness.
To address these challenges, we first propose a self-supervised pretraining task inspired by (Zhou et al., 2023b) to capture geometric information like edge lengths and triangle areas and enhance the vertex representations of Att-GCRN, which is referred to as Att-GCRN-GI. Second, we introduce an edge-covering-oriented graph sampling method that models graph sampling as a set-covering problem (Young, 2008), aiming to minimize the number of subgraphs and sampling time while ensuring comprehensive edge coverage and then get a sparse heatmap graph directly from the sub-heatmaps. We next detail the construction of the sparse heatmap graph with Att-GCRN-GI in the following, while leaving the architectural details of Att-GCRN-GI to Appendix A.2, as it is not the primary focus of this work.
Specifically, for each vertex $v_{i}$ of $\mathbb{G}$, we first get its corresponding subgraph $\mathbb{G}_{i}=(\mathcal{N}(v_{i},p),\mathcal{E}(\mathcal{N}(v_{i},p)))$, where $\mathcal{N}(v_{i},p)$ is to get $v_{i}$’s $p$ nearest neighbors and $\mathcal{E}(S)=\{(u,v)\in\mathbb{E}|u,v\in S\}$. We regard $C_{i}=\mathcal{E}(\mathcal{N}(v_{i},p))$ as a candidate set and can obtain subgraphs by identifying candidate sets that cover all edges for $n$ sets, thereby addressing a set-covering problem. Next, we apply a greedy algorithm (Young, 2008) to solve it, which iteratively selects the set containing the largest number of uncovered edges at each stage.
@@ 3.4 Hyper Tour Generation and Initialization
Motivated by the idea of “clustering first, route second” (Gillett and Miller, 1974), we propose addressing large-scale TSP instances by dividing them into clusters based on sparse heatmap graphs, and then pursuing a hyper tour among these clusters. Complex optimal solutions for large-scale TSP instances often exhibit smaller, well-organized local structures, such as subloops or clusters. These local structures are crucial to the overall solution, as they capture the optimal arrangement and connections of adjacent or related elements. Heatmaps can help identify these local structures by highlighting edges with high local probabilities, facilitating the grouping of vertices within the same sub-tour into a cluster. Each cluster is then abstracted into a supernode, a single vertex representing all vertices within the cluster—allowing us to form a smaller TSP instance while preserving the essential structure of the original graph. We solve this reduced problem and refer to its solution as the “hyper tour”, which determines how the supernodes are connected and serves as a guide for generating and optimizing the initial solution.
@@ 3.5 Targeted Neighborhood Search
Table 1: Performance on generated datasets (Luo et al., 2025). “N/A” indicates failure to solve due to memory constraints.
@@
Concorde (Cook et al., 2011) 23.12 (0%) 5.78h N/A N/A N/A N/A N/A N/A N/A N/A
LKH3 (Helsgaun, 2017) 23.12 (0%) 1.69h 50.90 (0%) 2.66h 71.77 (0%) 2.52h 101.31 (0%) 7.33h 159.93 (0%) 50.3h
AM-greedy (Kool et al., 2019) 33.55 (45.1%) 3.45m 96.74 (90.1%) 4.03m 153.42 (113.8%) 2.13m 244.20 (141.0%) 5.13m N/A N/A
GCN-BS (Joshi et al., 2019) 51.1 (121.0%) 3.23h N/A N/A N/A N/A N/A N/A N/A N/A
AttGCRN+MCTS (Fu et al., 2021)
@@
DIMES (Qiu et al., 2022) 23.69 (2.5%) 4.62h 52.37 (2.9%) 3.47h 74.06 (3.2%) 3.57h N/A N/A N/A N/A
DIFUSCO (Sun and Yang, 2023) 23.39 (1.2%) 0.41h 52.04 (2.2%) 0.71h 73.62 (2.6%) 0.79h N/A N/A N/A N/A
SoftDist (Xia et al., 2024) 23.63 (2.2%) 9.88m 52.25 (2.7%) 1.03h 74.03 (3.1%) 1.05h N/A N/A N/A N/A
SO (Cheng et al., 2023) 23.77 (2.8%) 0.93h 52.60 (3.3%) 1.92h 74.30 (3.5%) 2.03h 105.03 (3.7%) 8.92h N/A N/A
UDC (Zheng et al., 2024) 23.53(1.8%) 0.57h 33.26(2.5%) 0.21h N/A N/A N/A N/A N/A N/A
DRHG (Li et al., 2025a) 23.19 (0.3%) 2.31h 51.39 (0.9%) 3.13h 72.85 (1.3%) 4.26h 103.75 (2.4%) 6.14h 167.23 (4.6%) 9.35h
INViT (Fang et al., 2024) 24.50 (6.0%) 0.26h 54.19 (6.5%) 0.76h 76.12 (6.1%) 0.73h 108.49 (7.1%) 5.68h 171.38 (7.2%) 16.76h
PRC (Luo et al., 2025) 23.37 (1.1%) 3.47h 51.67 (1.3%) 6.23h 73.08 (1.8%) 3.71h 103.66 (2.3%) 5.93h 163.95 (2.5%) 16.92h
DualOpt (Zhou et al., 2025) 23.31 (0.8%) 0.31h 51.56 (1.3%) 0.64h 72.62 (1.2%) 0.87h 102.90 (1.6%) 2.28h 162.81 (1.8%) 7.83h